# An MLIR-based Compilation Framework for Control Flow Management on CGRAs

### v1

@@ 4. Compilation framework
Figure 6 illustrates a bird’s eye view of the compilation process, from high-level C/C++ code to low-level assembly generation. We first rely on the MLIR frontend to perform code parsing and generate the CDFG abstraction at a high level, independent of any specific backend. Section 5.1 introduces platform-independent optimizations, aiming to reduce control overhead without relying on specific hardware requirements. Sections 5.2 and 5.3 describe hardware compatibility transformations, which are applied before any mapping-based methods to address architecture-specific differences. As a lower-level optimization, modulo scheduling of loop is instead discussed in Section 6. The modulo scheduling results, which potentially map large subgraphs of target applications, may severely constraints the overall mapping. In Section 7, we explain how we address this lack of flexibility in our framework, seamlessly integrating modulo-scheduled loops in CDFG compilation. Finally, section 7.5 discusses the register allocation process, which is performed as a final step in our methodology.
Figure 6. Birds-eye view of the proposed CDFG compilation flow, enabling the deployment of kernel function on a CGRA mesues. The framework comprises a MLIR-based pre-processing front-end, back-end passes, ILP-based mapping, CFG transformations to support modulo scheduling and register allocation.
@@ 6. IR transformation for modulo scheduling
Figure 8. CFG transformation to enable loop-level parallelism on both IR level and assembly level. (a) original CFG with a single loop; (b) mapping solution for loop length $L$=4; (c) Original CFG for loop iteration; (d) original execution cycles; (e) transformed CFG with three blocks. The newly transformed loop block enables the simultaneous execution of two iterations in parallel; (f) New loop block has a reduced loop length $L^{\prime}=II=2$, while the original DFG is computed in two sequential iterations; (g) New CFG graph with new loop kernel encapsulates multiple iterations; (h) efficient execution cycles.
Input: loop length $L$; initial interval $II$;
@@
for $l^{\prime}$=$\lceil L/II\rceil-1:1$ do
@@ 7. CDFG mapping
Figure 9(a) illustrates the scheduling methodology for mapping a CDFG, which is decomposed into DFG scheduling via liveness analysis, as detailed in Section 7.1. Inter-BB live values are stored in physical registers, with strategies for managing them discussed in Section 7.2. Each DFG is mapped using the ILP model described in Section 7.3. When the ILP model successfully generates a solution, the mapping results for values propagated across BBs are recorded for CDFG-level integration (highlighted in the yellow table in Figure 9(a)). Subsequent DFGs that consume these values must adhere to the established mappings.
If the ILP model fails to find a feasible solution, a failure handler transforms the DFG by additional routing and DFG splitting. Details of this process are provided in Section 7.4. If the DFG is split in a node that is evicted to memory, the evicted values and their corresponding storage locations are recorded (highlighted in the grey table in Figure 9(a)), and any future references to these values are replaced with load operations from the specified memory locations. Finally, Section 7.5 describes the integration of all basic block mappings into the final assembly code generation.
@@ 7.1. Decomposition of CDFG to individual DFGs
A CDFG consists of connected basic blocks, allowing values to propagate across them. The interface between basic blocks and the CDFG is managed through live-in and live-out values, which define how data flows between blocks. A value is considered as live-in for a basic block $b$ if it is either used before being defined within the block or originates from a predecessor. This relationship is expressed as:
@@
These equations govern value propagation across basic blocks, ensuring that data dependencies are respected by tracking the live-in and live-out values. Algorithm 2 shows this process, which involves iteratively calculating which variables are live at different program points by analyzing the CFG and variable usage patterns (compiler_design, ).
The computed live values indicate that they are propagated across basic blocks. Their producer PEs are initially marked as unknown, while their register index is indicated either as ‘E’xternal or ‘Internal’, according to the heuristic described in the following section (see Figure 9(a)). ‘E’ values are stored in the output register (Rout) of PEs, allowing them to be read by neighboring cells but preventing the PE from executing other operations until the value has been completely consumed. ‘I’ values are stored in a PE register file (RF), which does not block computation but makes the value only accessible locally to a PE, as shown in Figure 2. Once the basic block mapper (described in Section 7.3) determines the production PE from their defining block, the PE index in the table is updated to ensure that consumer(s) operation of each value can access them.
@@ 7.2. Value assignment across basic blocks
where $\theta$ is a user-defined threshold, and $j$ is the index of its consumer. This strategy prevents a PE from being blocked by long dependency paths while ensuring the value remain accessible to all connected PEs.
@@ 7.3. Basic block mapper
Global live-in constraints: The consumer operations must adhere to specific spatial restrictions to utilize the live-ins generated outside the blocks, which are placed either in internal or external registers. If a livein variable $v_{k}^{\text{E}}$ is located in PE $p$ externally, then the execution units for all consumer $p_{I_{j}}$ where $I_{j}\in Cons(v_{k})$ should be the neighbor or $p$ itself. If the CGRA present nearest-neighbor connections, the constraint is formulated as:
(4) $p_{I_{j}}=$ $\text{T}(p)\ ||\text{B}(p)\ ||\text{L}(p)\ ||\text{R}(p\ )||p~,\forall I_{j}\in Cons(v_{k}^{\text{E}})$
@@
(9) $p_{I_{m}}=\text{T}(p_{I_{n}})\ ||\text{B}(p_{I_{n}})\ ||\text{L}(p_{I_{n}})\ ||\text{R}(p_{I_{n}}\ )||p_{I_{n}}~,\forall j\in Prod(I_{n})\vskip-3.00003pt$
@@ 7.4. IR transformation to handle ILP failure
The ILP model does not guarantee the existence of a solution due to CGRA hardware constraints. We herein discuss a strategy that counters two key sources of unfeasible solutions: routing limitations and register overflow. Figure 10 illustrates the IR of a block facing both failure scenarios. Figure 10(b) shows the pre-scheduling operations with live value placements, while the first row of Figure 10(a) indicates their locations. Value %4 encounters routing limitations as %0 is only accessible by PE0, while PE0 cannot reach PE3 for %1. This conflict is resolved by inserting a move operation (add %1, 0) to route %1 from PE3 to PE1. Value %5 cannot use %2 and %3, as they are stored in different PEs’ internal registers. PE3 is blocked from popping %3 until its external register value is consumed. To address this, %3 is evicted to memory, and a load operation is inserted. Figure 10(c) shows the transformed IR with the additional move and load operations, enabling a feasible mapping solution.
@@
If a value $v$ leads to the ILP model failure is used and defined only inside the basic block $i$, $v\notin use[b_{j}]~\forall j$, resulting $v\notin In\left[b_{j}\right]~\forall j$. Moreover, $v\notin In\left[b_{i}\right]$ because $v\notin use\left[b\right]-def\left[b\right]$ according to Equation 1. The value is in this case not recorded in the live value across BBs and cannot affect the other basic blocks.
@@
In case a load and store are inserted for $v$, let’s define its consumer as $b_{j}$ and its producer block $b_{i}$. For the producer block, the define and use sets remain the same: $use\{b_{i}\}^{\prime}=use\{b_{i}\}$, and $def\{b_{i}\}^{\prime}=def\{b_{i}\}$. For the consumer block, $use\{b_{j}\}^{\prime}=use\{b_{i}\}-v$, and $def\{b_{i}\}^{\prime}=def\{b_{i}\}+v_{0}$, where $v_{0}$ is the loaded value to replace $v$. This transformation updates the value across BBs by removing $v$ from BB propagations according to Equation 1. Figure 10(c) shows evicting value % 3 from the register to the memory. This still abides to the scheduler results, as it represents a relaxation of the constraints in Section 7.3, meaning the previous scheduling remains valid except a store operation need to be inserted after the produce of the split value %3. This process involves identifying an available neighboring PE after the split value has been produced. If no available PE exists, additional cycles are introduced to accommodate the store operation, without the need for re-scheduling.
@@ 7.6. Back end
As a final step for CDFG mapping, register allocation is performed. Operands from neighbor PEs are retrieved from their output registers, eliminating the need for register allocation in local RF when data is routed from one PE to another. Internal register allocation is performed per-PE, similarly to state of the art approaches such as (hack2006register, ). This step involves constructing a liveness graph for each PE and applying graph coloring to assign distinct physical registers to values that interfere with one another. After register allocation, a final rewriting pass converts the MLIR code to the needed input for the CGRA assembler. During the MLIR-to-CGRA assembler conversion, redundant jumps to basic blocks placed immediately afterwards in the instruction placements are removed. The assembler is then translated into binary form for execution on the target hardware.
@@ 8. Experimental framework
SoA methods. We compare our method and target platform with state-of-the-art mapping methods and their respective hardware requirements in Table 1. EffiMap (pc_control1, ) supports CDFG mappings with PC-control, while bi-directional jumps are implemented by a dedicated control unit. Constant register files are utilized to generate and manage constant values. Conversely, we achieve these functionalities entirely through software techniques, as explained in Section 5.3. A bi-directional jump is replaced with a one-direction conditional jump, the false case is executed with an unconditional jump, and constants are generated through combinations of arithmetic operations. Marionette (control_plane, ) also supports CDFG but requires complex hardware components, including a reconfiguration unit and a CFG control network to propagate the DFG switch decision. SatMapIt (satjournal, ) only focuses on DFG mapping, leveraging external CPU for CFG management. We consider this methodology both as a stand-alone baseline strategy, and as the modulo scheduler employed in our framework.
@@ 9. Results
We evaluate our approach from multiple viewpoints. First, we showcase that our CDFG stance effectively increase the compilation scope with respect to DFG-based alternatives. Then, we report the effect of the devised CFG-based optimizations on compile time. Furthermore, we discuss the run-time performance improvements deriving from each optimization technique introduced in this work. Finally, we compare our compilation framework with state-of-the-art approaches and their respective target platforms, demonstrating that our method offers higher mapping efficiency while requiring minimal hardware for controlling execution.
@@ 9.1. IR scope analysis
Figure 11 illustrates the applicable scope of DFG mappers, where the blue sections represent operations for control-free loops and the orange sections indicate operations outside the innermost loops, for which control flow considerations are required. The numbers and ratios shown in the pie charts indicate the count of operations and their proportion relative to the total number of operations in each application. Results indicate that the scope and complexity of CGRA kernels can be effectively increased by considering control flow. The CFG of FIR cannot be simplified to include a control-free loop that is targetable by DFG mappers, which must be mapped by a CDFG compiler. Moreover, they hint that a DFG-only approach can result in high overheads for the orchestration of CGRA configuration and input/output data transfers, as discussed in Section 9.4. Such overheads are entirely waived in our proposed strategy, as a single configuration is required for supporting the acceleration of the entirety of each kernel.
@@ 9.2. Compilation time
The compilation time encompasses the whole compilation process required to map an application, including the front, middle, and back end for IR transformation, operation mapping, and register allocation for assembly generation. Figure 12 illustrates the total compilation time targeting 3×3 and 4×4 CGRAs. In the plain compilation flow, operation mapping is performed directly on the IR generated by the MLIR front end. In contrast, the optimized flow processes the IR through CFG simplification (Section 5.1) and adaptation with modulo scheduling (Section 6). CDFG mapping is decomposed in our methodology into the mapping of its constituent basic blocks. Therefore, the total mapping time is primarily influenced by two factors: the number of operations within individual basic blocks and the total number of basic blocks. Mapping complexity increases exponentially with the number of operations in a basic block, making blocks with large DFG sizes the primary contributors to overall mapping time. Additionally, the presence of multiple basic blocks with high mapping complexity further extends compilation time. For instance, the gsm benchmark contains six basic blocks, each with fewer than five operations, resulting in relatively fast compilation. In contrast, sha comprises eight basic blocks, each containing over 25 operations, which significantly increases mapping time.
Operation mapping dominates the overall compilation time, whereas the time spent on IR transformation and register allocation is negligible. This is exemplified by the FIR benchmark, whose CFG cannot be simplified by our approach. Hence, the observed increase in compilation time is attributed solely to IR transformation, which contributes only a few milliseconds. The optimized implementation exhibits longer compilation times compared to the plain version for the conv1d benchmark, with compilation overhead increasing by several seconds. In the case of conv1d, the control-free kernels contain small counts of operations with reduced dependencies, making the plain version easier to find a solution. The modulo scheduler performs more extensive exploration to identify initiation intervals that can exploit the available parallelism. Indeed, most optimized compilation flow achieves faster compilation than the plain methodology. This improvement stems from CFG simplification, which restructures the CFG to be amenable to modulo scheduling. The SAT-based modulo scheduler employed in our flow reduces mapping time for computation-intensive loop kernels.
@@ 9.3. Impact of optimizations on run-time
Three factors influence the runtime efficiency of the mapping results produced by the proposed framework, ranked in order of their impact: (1) the register attributes used to assign inter-basic-block values, as discussed in Section 7.2; (2) CFG simplification techniques that change the CFG structure, detailed in Section 5.1; and (3) transformations based on modulo scheduling to enable loop-level parallelism, described in Section 6. Figure 13 illustrates the performance improvements achieved through different combinations of optimizations. Plain and Optimized compilation settings are defined as in Section 9.2— in Plain no optimization is applied, while in Optimized all are employed. Additionally, we evaluate intermediate configurations that (1) apply only values allocations in both internal and external registers, and (2) apply CFG simplifications further.
@@
The placement of values propagated across basic blocks significantly impacts convolution and linear algebra benchmarks, as these workloads carry high data dependencies through nested loops for computing memory addresses. For a 4×4 CGRA running convolutional benchmarks, selectively placing partial variables in external registers expands the access range, resulting in an 6.6% acceleration gain on average. Combinational use of the CFG transformations simplifies the control flow, particularly in gsm, where five blocks are merged into a single self-iterating loop. This CFG simplification enables multiple blocks to be consolidated, allowing merged operations to be scheduled simultaneously. As a result, acceleration gains of 26.8% and 26.4% are achieved on 3×3 and 4×4 CGRAs, respectively. For conv1d on a 4×4 CGRA and gemm on a 3×3 CGRA, performance improvements reach up to 48% and 45%, respectively. Loops with lower nested levels such as gsm and gemm obtain 316% and 255% runtime faster on a 4x4 CGRA. On average, modulo scheduling folds the innermost loop, achieving 197% on 3x3 and 223% acceleration on 4x4 CGRAs, indicating $L\approx 2*II$ achieved by the modulo scheduler. CGRAs with larger grid sizes achieve higher speedups, as the presence of more hardware resources enables higher parallelism. In general, the introduced optimization are highly beneficial on all considered benchmarks. FIR is an outlier, as its CDFG structure presents a write operations in one branch inside an innermost loop, which disallow basic block merging.

### v2

@@ 4. Compilation framework
Figure 6 illustrates a bird’s eye view of the compilation process, from high-level C/C++ code to low-level assembly generation. We first rely on the MLIR frontend to perform code parsing and generate the CDFG abstraction at a high level, independent of any specific backend. Section 5.1 introduces platform-independent optimizations, with the aim of reducing control overhead without relying on specific hardware requirements. Sections 5.2 and 5.3 describe hardware compatibility transformations, which are applied before any mapping-based methods to address architecture-specific differences. As a lower-level optimization, modulo scheduling of loop is instead discussed in Section 6. The modulo scheduling results, which potentially map large subgraphs of target applications, may severely constrain the overall mapping. In Section 7, we explain how we address this lack of flexibility in our framework, seamlessly integrating modulo-scheduled loops in CDFG compilation. Finally, Section 7.5 discusses the register allocation process, which is performed as a final step in our methodology.
Figure 6. Birds-eye view of the proposed CDFG compilation flow, enabling the deployment of kernel function on a CGRA meshes. The framework comprises a MLIR-based pre-processing front-end, back-end passes, ILP-based mapping, CFG transformations to support modulo scheduling and register allocation.
@@ 6. IR transformation for modulo scheduling
Figure 8. CFG transformation to enable loop-level parallelism on both IR level and assembly level. (a) original CFG with a single loop; (b) mapping solution for loop length LL=4; (c) Original CFG for loop iteration; (d) original execution cycles; (e) transformed CFG with three blocks. The newly transformed loop block enables the simultaneous execution of two iterations in parallel; (f) New loop block has a reduced loop length $L^{\prime}=II=2$, while the original DFG is computed in two sequential iterations; (g) New CFG graph with new loop kernel encapsulates multiple iterations; (h) efficient execution cycles.
Input: loop length $L$; initial interval IIII;
@@
for $l^{\prime}$$\lceil L/II\rceil-1:1$ do
@@ 7. CDFG mapping
Figure 9(a) illustrates the scheduling methodology for mapping a CDFG, which is decomposed into DFG scheduling through liveness analysis, as detailed in Section 7.1. Inter-BB live values are stored in physical registers, with strategies for managing them discussed in Section 7.2. Each DFG is mapped using the ILP model described in Section 7.3. When the ILP model successfully generates a solution, the mapping results for values propagated across BBs are recorded for CDFG-level integration (highlighted in the yellow table in Figure 9(a)). Subsequent DFGs that consume these values must adhere to the established mappings.
If the ILP model fails to find a feasible solution, a failure handler transforms the DFG by adding routing and DFG splitting. Details of this process are provided in Section 7.4. If the DFG is split in a node that is evicted to memory, the evicted values and their corresponding storage locations are recorded (highlighted in the grey table in Figure 9(a)), and any future references to these values are replaced by load operations from the specified memory locations. Finally, Section 7.5 describes the integration of all basic block mappings into the generation of the final assembly code.
@@ 7.1. Decomposition of CDFG to individual DFGs
A CDFG consists of connected basic blocks, allowing values to propagate through them. The interface between basic blocks and the CDFG is managed through live-in and live-out values, which define how data flows between blocks. A value is considered as live-in for a basic block $b$ if it is used before being defined within the block or originates from a predecessor. This relationship is expressed as:
@@
These equations govern the propagation of values across basic blocks, ensuring that data dependencies are respected by tracking live-in and live-out values. Algorithm 2 shows this process, which involves iteratively calculating which variables are live at different program points by analyzing the CFG and variable usage patterns (compiler_design, ).
The computed live values indicate that they are propagated across basic blocks. Their producer PEs are initially marked as unknown, while their register index is indicated either as ‘External or ‘Internal’, according to the heuristic described in the following section (see Figure 9(a)). ‘E’ values are stored in the output register (Rout) of PEs, allowing them to be read by neighboring cells but preventing the PE from executing other operations until the value has been completely consumed. ‘I’ values are stored in a PE register file (RF), which does not block the computation, but makes the value only accessible locally to a PE, as shown in Figure 2. Once the basic block mapper (described in Section 7.3) determines the PE production from their defining block, the PE index in the table is updated to ensure that the consumer(s) operation of each value can access them.
@@ 7.2. Value assignment across basic blocks
where $\theta$ is a user-defined threshold, and $j$ is the index of its consumer. This strategy prevents a PE from being blocked by long dependency paths while ensuring the value remains accessible to all connected PEs.
@@ 7.3. Basic block mapper
Global live-in constraints: Consumer operations must adhere to specific spatial restrictions to utilize live-ins generated outside the blocks, which are placed in internal or external registers. If a livein variable $v_{k}^{\text{E}}$ is located in PE $p$ externally, then the execution units for all consumer $p_{I_{j}}$ where $I_{j}\in Cons(v_{k})$ should be the neighbor or $p$ itself. If the CGRA present nearest-neighbor connections, the constraint is formulated as:
[…]
@@
[…]
@@ 7.4. IR transformation to handle ILP failure
The ILP model does not guarantee the existence of a solution due to CGRA hardware constraints. We herein discuss a strategy that counters two key sources of unfeasible solutions: routing limitations and register overflow. Figure 10 illustrates the IR of a block facing both failure scenarios. Figure 10(b) shows the pre-scheduling operations with live value placements, while the first row of Figure 10(a) indicates their locations. Value %4 encounters routing limitations as %0 is only accessible by PE0, while PE0 cannot reach PE3 for %1. This conflict is resolved by inserting a move operation (add %1, 0) to route %1 from PE3 to PE1. Value %5 cannot use %2 and %3, as they are stored in different PEs’ internal registers. PE3 is blocked from popping %3 until its external register value is consumed. To address this, %3 is evicted to memory, and a load operation is inserted. Figure 10(c) shows the transformed IR with additional move and load operations, allowing a feasible mapping solution.
@@
If a value $v$ leads to the failure of the ILP model is used and defined only inside the basic block $i$, $v\notin use[b_{j}]~\forall j$, the result $v\notin In\left[b_{j}\right]~\forall j$. Moreover, $v\notin In\left[b_{i}\right]$ because $v\notin use\left[b\right]-def\left[b\right]$ according to Equation 1. In this case, the value is not recorded in the live value across BBs and cannot affect the other basic blocks.
@@
In case a load and store are inserted for $v$, let’s define its consumer as $b_{j}$ and its producer block $b_{i}$. For the producer block, the define and use sets remain the same: $use\{b_{i}\}^{\prime}=use\{b_{i}\}$, and $def\{b_{i}\}^{\prime}=def\{b_{i}\}$. For the consumer block, $use\{b_{j}\}^{\prime}=use\{b_{i}\}-v$, and $def\{b_{i}\}^{\prime}=def\{b_{i}\}+v_{0}$, where $v_{0}$ is the loaded value to replace $v$. This transformation updates the value across BBs by removing $v$ from BB propagations according to Equation 1. Figure 10(c) shows evicting value % 3 from the register to the memory. This still abides by the scheduler results, as it represents a relaxation of the constraints in Section 7.3, which means that the previous scheduling remains valid except that a store operation need to be inserted after the production of the split value %3. This process involves identifying an available neighboring PE after the split value has been produced. If no available PE exists, additional cycles are introduced to accommodate the store operation, without the need for re-scheduling.
@@ 7.6. Back end
As a final step for CDFG mapping, register allocation is performed. Operands from neighbor PEs are retrieved from their output registers, eliminating the need for register allocation in local RF when data is routed from one PE to another. Internal register allocation is performed per-PE, similarly to state-of-the-art approaches such as (hack2006register, ). This step involves constructing a liveness graph for each PE and applying graph coloring to assign distinct physical registers to values that interfere with one another. After register allocation, a final rewriting pass converts the MLIR code to the input needed for the CGRA assembler. During the MLIR-to-CGRA assembler conversion, redundant jumps to basic blocks placed immediately afterwards in the instruction placements are removed. The assembler is then translated into binary form for execution on the target hardware.
@@ 8. Experimental framework
SoA methods. We compare our method and target platform with state-of-the-art mapping methods and their respective hardware requirements in Table 1. EffiMap (pc_control1, ) supports CDFG mappings with PC-control, while bi-directional jumps are implemented by a dedicated control unit. Constant register files are utilized to generate and manage constant values. Conversely, we achieve these functionalities entirely through software techniques, as explained in Section 5.3. A bi-directional jump is replaced with a one-direction conditional jump, the false case is executed with an unconditional jump, and constants are generated through combinations of arithmetic operations. Marionette (control_plane, ) also supports CDFG but requires complex hardware components, including a reconfiguration unit and a CFG control network to propagate the DFG switch decision. SatMapIt (satjournal, ) focuses only on DFG mapping, leveraging external CPUs for CFG management. We consider this methodology both as a stand-alone baseline strategy and as the modulo scheduler employed in our framework.
@@ 9. Results
We evaluate our approach from multiple points of view. First, we demonstrate that our CDFG stance effectively increases the compilation scope with respect to DFG-based alternatives. Then, we report the effect of the devised CFG-based optimizations on compile time. Furthermore, we discuss the run-time performance improvements that result from each optimization technique introduced in this work. Finally, we compare our compilation framework with state-of-the-art approaches and their respective target platforms, demonstrating that our method offers higher mapping efficiency while requiring minimal hardware for controlling execution.
@@ 9.1. IR scope analysis
Figure 11 illustrates the applicable scope of DFG mappers, where the blue sections represent operations for control-free loops and the orange sections indicate operations outside the innermost loops, for which control flow considerations are required. The numbers and ratios shown in the pie charts indicate the count of operations and their proportion relative to the total number of operations in each application. The results indicate that the scope and complexity of the CGRA kernels can be effectively increased by considering the control flow. The CFG of FIR cannot be simplified to include a control-free loop that is targetable by DFG mappers, which must be mapped by a CDFG compiler. Moreover, they hint that a DFG-only approach can result in high overheads for the orchestration of CGRA configuration and input/output data transfers, as discussed in Section 9.4. Such overheads are entirely waived in our proposed strategy, as a single configuration is required to support the acceleration of the entirety of each kernel.
@@ 9.2. Compilation time
The compilation time encompasses the whole compilation process required to map an application, including the front, middle, and back end for IR transformation, operation mapping, and register allocation for assembly generation. Figure 12 illustrates the total compilation time targeting 3×3 and 4×4 CGRAs. In the plain compilation flow, the operation mapping is performed directly on the IR generated by the MLIR front end. In contrast, the optimized flow processes the IR through CFG simplification (Section 5.1) and adaptation with modulo scheduling (Section 6). CDFG mapping is decomposed into the mapping of its basic constituent blocks in our methodology. Therefore, the total mapping time is mainly influenced by two factors: the number of operations within individual basic blocks and the total number of basic blocks. Mapping complexity increases exponentially with the number of operations in a basic block, making blocks with large DFG sizes the primary contributors to overall mapping time. Additionally, the presence of multiple basic blocks with high mapping complexity further extends the compilation time. For example, the gsm benchmark contains six basic blocks, each with fewer than five operations, resulting in a relatively fast compilation. In contrast, sha comprises eight basic blocks, each containing more than 25 operations, significantly increasing the mapping time.
Operation mapping dominates the overall compilation time, whereas the time spent on IR transformation and register allocation is negligible. This is exemplified by the FIR benchmark, whose CFG cannot be simplified by our approach. Hence, the observed increase in compilation time is attributed solely to IR transformation, which contributes only a few milliseconds. The optimized implementation exhibits longer compilation times compared to the plain version for the conv1d benchmark, with compilation overhead increasing by several seconds. In the case of conv1d, the control-free kernels contain small counts of operations with reduced dependencies, making the plain version easier to find a solution. The modulo scheduler performs more extensive exploration to identify initiation intervals that can exploit the available parallelism. In fact, most optimized compilation flow achieves faster compilation than the plain methodology. This improvement comes from CFG simplification, which restructures the CFG to be amenable to modulo scheduling. The SAT-based modulo scheduler employed in our flow reduces mapping time for computation-intensive loop kernels.
@@ 9.3. Impact of optimizations on run-time
Three factors influence the runtime efficiency of the mapping results produced by the proposed framework, ranked in order of their impact: (1) the register attributes used to assign inter-basic-block values, as discussed in Section 7.2; (2) CFG simplification techniques that change the CFG structure, detailed in Section 5.1; and (3) transformations based on modulo scheduling to enable loop-level parallelism, described in Section 6. Figure 13 illustrates the performance improvements achieved through different combinations of optimizations. Plain and Optimized compilation settings are defined as in Section 9.2— in Plain no optimization is applied, while in Optimized all are employed. Additionally, we evaluate intermediate configurations that (1) apply only value allocations in both internal and external registers, and (2) apply CFG simplifications further.
@@
The placement of values propagated across basic blocks significantly impacts convolution and linear algebra benchmarks, as these workloads carry high data dependencies through nested loops for computing memory addresses. For a 4×4 CGRA running convolutional benchmarks, selectively placing partial variables in external registers expands the access range, resulting in an 6.6% acceleration gain on average. Combinational use of CFG transformations simplifies control flow, particularly in gsm, where five blocks are merged into a single self-iterating loop. This CFG simplification enables multiple blocks to be consolidated, allowing merged operations to be scheduled simultaneously. As a result, acceleration gains of 26.8% and 26.4% are achieved on 3×3 and 4×4 CGRAs, respectively. For conv1d on a 4×4 CGRA and gemm on a 3×3 CGRA, performance improvements reach up to 48% and 45%, respectively. Loops with lower nested levels such as gsm and gemm obtain 316% and 255% runtime faster on a 4x4 CGRA. On average, modulo scheduling folds the innermost loop, achieving 197% on 3x3 and 223% acceleration on 4x4 CGRAs, indicating $L\approx 2*II$ achieved by the modulo scheduler. CGRAs with larger grid sizes achieve higher speedups, as the presence of more hardware resources enables greater parallelism. In general, the introduced optimization is highly beneficial on all considered benchmarks. FIR is an outlier, as its CDFG structure presents a write operations in one branch inside an innermost loop, which disallow basic block merging.